OpenStudy (anonymous):

3. Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. 3x – 4y > 0 x – 5y > 0 A. (6, 0) B. (6, 2) C. (–4, –2) D. (0, 0)

OpenStudy (anonymous):

thanks math male

OpenStudy (anonymous):

Graph it.

OpenStudy (anonymous):

and i also have those 2 screen shots

OpenStudy (mathmale):

Hello! I'd suggest that you start by temporarily replacing those inequality symbols by = signs. First: Are you able to graph linear equations? Could you graph 3x - 4y = 0?

OpenStudy (anonymous):

okay let me see thanks for the help btw

OpenStudy (anonymous):

First figure out the inequalities then graph it on to a graph.

OpenStudy (anonymous):

Do you know how to figure the inequality out and graph coordinates?

OpenStudy (mathmale):

In this particular case, I'd suggest that you solve 3x - 4y = 0 for y, so that you'll have the slope of this straight line. Can you do that?

OpenStudy (anonymous):

honestly I'm really bad at math so I'm not very good at graphs either

OpenStudy (anonymous):

yes i think i can

OpenStudy (anonymous):

maybe

OpenStudy (mathmale):

@Anonymouseee : While you are welcome to post here, please read other posts to ensure that you're not repeating the very same information.

OpenStudy (mathmale):

Please solve 3x - 4y = 0 for y.

OpenStudy (anonymous):

can you please look at the screen shots?

OpenStudy (anonymous):

okay i will

OpenStudy (anonymous):

Tell me if you can or not. If you know how to figure out an inequality by your self and put the coordinates on the graph by yourself without help, I can just give you an answer.

OpenStudy (anonymous):

@mathmale would it be d?

OpenStudy (anonymous):

yea its d

OpenStudy (anonymous):

(0,0)?

OpenStudy (anonymous):

okay thanks what about the rest can you help me out? @Anonymouseee

OpenStudy (anonymous):

and mathmale

OpenStudy (anonymous):

If you graph it and draw it out the line goes through 0,0

OpenStudy (anonymous):

you guys are a big help thanks

OpenStudy (anonymous):

both lines***

OpenStudy (anonymous):

and my apologies mathmale.

OpenStudy (anonymous):

can you guys check out the screen shots @Anonymouseee @mathmale

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Actually post another thread for the other questions, since this thread has been completed.

OpenStudy (anonymous):

i think its a or b

OpenStudy (anonymous):

okay

OpenStudy (mathmale):

My understanding of your initial question is that you have the inequalities 3x – 4y > 0 and x – 5y > 0 Please double-check: Are these inequalities the same as in your three screen shots? If so, fine. If not, then the screen shots are irrelevant.

OpenStudy (anonymous):

@mathmale im not really sure what you are saying sorry. I'm not very good in anything in school what so ever

OpenStudy (anonymous):

for me double checking is hard

OpenStudy (mathmale):

I'd really suggest that you graph 3x-4y=0 and then x-5y=0. Then we will graph the related inequalities 3x-4y=0 and x-5y=0.

OpenStudy (anonymous):

for which question

OpenStudy (anonymous):

oh

OpenStudy (mathmale):

the one you originally posted.

OpenStudy (anonymous):

the non screen shot?

OpenStudy (anonymous):

isn't the one i originally posted 0,0

OpenStudy (mathmale):

Right. I just don't see the relevance of the screen shots. Please let's focus on one problem at a time. Let's get started: If 3x-4y=0, then -4y=-3x. Divde both terms by -4. What do you get?

OpenStudy (anonymous):

you get 0 0 right?

OpenStudy (anonymous):

i attempted it and i got 0,0 and my computer is on 11% so i have to hurry up

OpenStudy (mathmale):

In that case it might be better were we to continue in the morning. Where are you located (which state)?

OpenStudy (anonymous):

Charge it.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

california

OpenStudy (anonymous):

its 10:50

OpenStudy (anonymous):

pm

OpenStudy (anonymous):

and school starts at 7:45 am

OpenStudy (anonymous):

can @Anonymouseee @mathmale help me out tomorrow i might not need it but will you guys be on 2morrow?

OpenStudy (anonymous):

I stay up till like 2 am studying and wake up at 6 am haha.

OpenStudy (anonymous):

haha

OpenStudy (mathmale):

I'm in CA also. If you're up early enough I'd be glad to work with you, we'd need only 15-20 minutes for this problem.

OpenStudy (anonymous):

Well I am busy in the mornings so let mathmale handle it.

OpenStudy (anonymous):

okay thanks

OpenStudy (mathmale):

Again, @poptartmn , what time do you expect to be up tomorrow, and do you think you'll have time to work on openstudy before school?

OpenStudy (anonymous):

Yep. Hope you get figure it out man.

OpenStudy (anonymous):

i don't think i need it but I'm a fan of both of you

OpenStudy (anonymous):

my big sister just helped me out because she was getting mad that the computer is dying

OpenStudy (anonymous):

@mathmale no i don't think so and my big sister helped me

OpenStudy (anonymous):

but thanks for helping me out :)

OpenStudy (mathmale):

OK. Hope to work with you another time. Good night.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

can u help me with last ?

OpenStudy (anonymous):

computer is on 7 % and my big sister needs to use it i will just do it later thanks for the help tho